# Statistical Taylor Expansion111

### v9

@@ 2.2 Distributional Zero and Distributional Pole
figure The probability density function for $(x\pm 1)^{2}$, for different $x$ as shown in the legend. The $y$-axis is scaled as $\sqrt{\tilde{y}}$.
@@
figure The probability density function for $\tilde{y}=\sqrt{x\pm 1}$, for different $x$ as shown in the legend. The $x$-axis is scaled as $\tilde{y}^{2}$.
Let $\rho(\tilde{x},x,\delta x)$ be the distribution density function of random variable $\tilde{x}$ with mean $x$ and deviation $\delta x$. Let $\rho(\tilde{z})$ be the normalized form of $\rho(\tilde{x},x,\delta x)$, in which $\tilde{z}\equiv\frac{\tilde{x}-x}{\delta x}$, e.g., Normal distribution $N(\tilde{z})$ is the normalized form of Gaussian distribution.
Let $\tilde{y}=f(\tilde{x})$ be a strictly monotonic function, so that $\tilde{x}=f^{-1}(\tilde{y})$ exists. Formula (2.2) is the distribution density function of $\tilde{y}$ [2][5]. In Formula (2.2), the same distribution can be expressed in either $\tilde{x}$ or $\tilde{y}$ or $\tilde{z}$ , which are different representations of the same underlying random variable. Using Formula (2.2), Formula (2.3) gives the $\rho(\tilde{y},y,\delta y)$ for $x^{c}$ if $\rho(\tilde{x},x,\delta x)$ is Gaussian.
$\rho(\tilde{z})d\tilde{z}=$ $\rho(\tilde{x},x,\delta x)d\tilde{x}=\rho(f^{-1}(\tilde{y}),x,\delta x)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},y,\delta y)d\tilde{y};$ (2.2)
$y=x^{c}:\;\;\;\rho(\tilde{y},y,\delta y)=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\delta x}N(\frac{\tilde{y}^{\frac{1}{c}}-x}{\delta x});$ (2.3)
Viewed in the $f^{-1}(\tilde{y})$ coordinate, $\rho(\tilde{y},y,\delta y)$ is $\rho(\tilde{x},x,\delta x)$ modulated by $\frac{d\tilde{x}}{d\tilde{y}}=1/f^{(1)}_{x}$, in which $f^{(1)}_{x}$ is the first derivative of $f(x)$ to $x$. A distributional zero of the uncertainty distribution happens when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},y,\delta y)=0$, while a distributional pole happens when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},y,\delta y)=\infty$. Zeros and poles provides the strongest local modulation to $\rho(\tilde{x},x,\delta x)$:
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the result distribution is same as the original distribution, as $\rho(\tilde{y},y,\delta y)=\rho(\tilde{y},\alpha+\beta x,\beta\delta x)$ [5]. A linear transformation generates neither distributional zero nor distributional pole according to Formula (2.2).
@@
In both Figure 2.2 and 2.2, the result probability density function in the $\tilde{z}$ representation resembles the input distribution more when the mode of the distributions is further away from either distributional pole or distributional zero.
@@ 2.3 Statistical Taylor Expansion
Formula (2.2) gives the uncertainty distribution of an analytic function. However, normal scientific and engineering calculations usually do not care about the result distribution, but just few statistics of the result, such as the result mean and deviation [2][3]. Such simple statistics is the result of statistical Taylor expansion. In $x\pm\delta x$, $\delta x$ can either be the sample deviation of the input, or the calculated result, so it is defined as uncertainty.
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines bounded momentum $\zeta(n,\kappa)$, in which $0<\varrho,\kappa$ are bounding factors (As described later, $\kappa$ determines $\varrho$ through the constraint $\zeta(0,\kappa)=0$). Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@
$f(x+\tilde{x})=$ $f(x+\tilde{z}\delta x)=\sum_{n=0}^{\infty}\frac{f^{(n)}_{x}}{n!}\tilde{z}^{n}(\delta x)^{n};$ (2.5)
@@
$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.7)
@@
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{m}(\delta y)^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
$\delta^{2}f(x,y)=$ $\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
@@
2.4 Bounding Factor and Sample Count
[…]
figure The measured bounding leakage (as shown by the y-axis) for the different measuring bounding factor $\kappa_{s}$ (as shown by the x-axis) and sample count $N$ (as shown by the legend).
[…]
figure The measured bounding factor $\kappa$ (as shown by the y-axis on the left) and the corresponding measured bounding leakage $\epsilon(\kappa)$ (as shown by the y-axis on the right) for the different sample count $N$ (as shown by the x-axis) when the underlying distribution is either uniform or normal (as shown by the legend, with different measuring bounding factor for the normal distribution).
[…]
figure The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (as shown by the y-axis on the left) and the bounding leakage (as shown by the y-axis on the right) for the different bounding factor $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is the corresponding variance for $\kappa=5$.
[…]
figure The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for the different bounding factor $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), in which $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ so that $\delta^{2}x=(\delta x)^{2}$.
When the bounding factors $\varrho$ and $\kappa$ are limited, Formula (2.4) provides statistical bounding $x\in[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$. Statistical bounding is necessary for statistical Taylor expansion to converges in many cases, such as for $x^{c}$ and $\log(x)$. The probability $\epsilon$ for $x\not\in[\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$ is defined as the bounding leakage.
When sampling from a underlying distribution with the mean $\mu$ and deviation $\sigma$, the sample mean $\overline{x}$ and sample deviation $\delta x$ approaches $\mu$ and $\sigma$ respectively with the sample count $N$ as $\sqrt{N}$, so that the sample range $[\overline{x}-\kappa\delta x,\overline{x}+\kappa\delta x]$ is different from the actual range $[\mu-\kappa\sigma,\mu+\kappa\sigma]$, to have a different bounding leakage $\epsilon(\kappa,N)$ vs. $\epsilon(\kappa)$.
$\epsilon(\kappa)=$  $\;1-\frac{1}{2}\xi(\frac{\varrho}{\sqrt{2}})-\frac{1}{2}\xi(\frac{\kappa}{\sqrt{2}});$ (2.11)
$\epsilon(\kappa,N)=$  $\;1-\frac{1}{2}\xi(\frac{|\varrho\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa\delta x+\overline{x}|}{\sqrt{2}});$ (2.12)
[…]
Generally, $\kappa$ in Formula (2.4) increases with the sample count $N$, which has to be sufficiently large to get accurate result using statistical Taylor expansion. Let $\widehat{\delta^{2}f}$ be the ideal variance when the result variance becomes stable with increasing $\kappa$. Figure 2.4 shows that $\delta^{2}f$ increases with $\kappa$ until reaches $\widehat{\delta^{2}f}$ for the selected functions. The result bounding factor is defined by Formula (2.13) using the result variance ratio, e.g., Formula (2.14) shows that the result $\kappa$ for $x\pm y$ is a variance weighted average of the input $\kappa(N_{x})$ and $\kappa(N_{y})$ through $\zeta(2,\kappa)$. Figure 2.4 also shows that $\kappa$ remains roughly a constant during calculation, with all the result variance ratios very close to that of $x$ for all the input $\kappa$.
$\zeta(2,\kappa)\equiv$ $\;\frac{\delta^{2}f}{\widehat{\delta^{2}f}};$ (2.13)
$x\pm y:$ $\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa(N_{x}))(\delta x)^{2}+\zeta(2,\kappa(N_{y}))(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.14)
The cutoff $\kappa$ for $\widehat{\delta^{2}f}$ is somewhat subjective. In this paper, it is when $\kappa$ reaches $5$ or the input sample count $N$ reaches $1000$ for Gaussian distribution.
As stated later, when $1/5<P(x)$, $\log(x\pm\delta x)$ or $(x\pm\delta x)^{c}$ diverges for $\widehat{\delta^{2}f}$. To approximate $\widehat{\delta^{2}f}$, an artificial $\kappa_{a}$ satisfying $P(x)^{2}\leq\zeta(2,\kappa_{a})/\kappa_{a}^{2}$ can be used to normalize the input variance $(\delta x)^{2}$ as $(\delta x)^{2}/\zeta(2,\kappa_{a})$, to achieve $\delta^{2}x=(\delta x)^{2}$ perfectly for all $\kappa$. Figure 2.4 shows that the result variances are close to those of $5<\kappa$, although as a first order approximation, the results depend on the actual $f(x)$. Figure 2.4 shows a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions disappears.
@@ 2.5 Bounding Symmetry and Asymptote
A bounding $[x-\varrho\delta x,x+\kappa\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. To achieve mean-preserving, $\kappa$ determines $\varrho$. With mean-preserving bounding, Formula (2.15) and (2.16) give the result for $x\pm y$, while Formula (2.17) and (2.18) give the result for $xy$:
@@
$\delta^{2}(x\pm y)$ $=\zeta_{x}(2)(\delta x)^{2}+\zeta_{y}(2)(\delta y)^{2};$ (2.16)
@@
$\delta^{2}(xy)$ $=\zeta_{x}(2)(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2)(\delta y)^{2}+\zeta_{x}(2)(\delta x)^{2}\zeta_{y}(2)(\delta y)^{2};$ (2.18)
@@ 2.5.1 Uniform Input Uncertainty
For uniform distribution, the bounding factor is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.19) gives $\zeta(2n)$ while $\zeta(2n+1)=0$.
@@ 2.5.2 Gaussian Input Uncertainty
If $\kappa=\infty$, $\zeta(2n)=(2n-1)!!$!, which causes Formula (2.37) for $\log(x\pm\delta x)$ and Formula (2.40) for $(x\pm\delta x)^{c}$ to diverge. When $\kappa$ has an upper bound, Formula (2.4) becomes Formula (2.20):
@@
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$! according to Formula (2.22). When $\kappa=5,n<5$, $\left|1-\eta(2n)/(2n-1)!!\right|<10^{-3}$.
@@ 2.5.3 An Input Uncertainty with Limited Range
Formula (2.24) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$. Formula (2.25) shows its mean-preserving equation. Formula (2.26) shows its bounded momentum. Formula (2.27) shows its asymptote. Formula (2.28) shows its bounding leakage.
$\rho(\tilde{x},x,\delta x)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;x=2\lambda;\;\;\;\delta^{2}x=2\lambda^{2};$ (2.24)
@@ 2.6 One Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(2n)\frac{1}{n!};$ (2.30)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.31)
@@
$\overline{\sin(x)}=$ $\sin(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.33)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.34)
@@
The result variance in statistical Taylor expansion shows the nature of the calculation, such as $\delta x\rightarrow P(e^{x})$, $\delta x\rightarrow\delta\sin(x)$, $P(x)\rightarrow P(x^{c})$, and $P(x)\rightarrow\delta\log(x)$.

### v10

@@ 2.2 Distributional Zero and Distributional Pole
figure The probability density function for $\tilde{y}=\tilde{x}^{2}$, for different $\mu$ as shown in the legend. The $\tilde{x}$ is Gaussian distributed with the distributional mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\sqrt{\tilde{y}}$.
@@
figure The probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, for different $\mu$ as shown in the legend. The $\tilde{x}$ is Gaussian distributed with the distributional mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\tilde{y}^{2}$.
Let $\rho(\tilde{x},\mu,\sigma)$ be the probability density function of random variable $\tilde{x}$ with the distributional mean $\mu$ and deviation $\sigma$. Let $\rho(\tilde{z})$ be the normalized form of $\rho(\tilde{x},\mu,\sigma)$, in which $\tilde{z}\equiv(\tilde{x}-\mu)/\sigma$, e.g., Normal distribution $N(\tilde{z})$ is the normalized form of Gaussian distribution.
Let $\tilde{y}=f(\tilde{x})$ be a strictly monotonic function, so that $\tilde{x}=f^{-1}(\tilde{y})$ exists. Formula (2.2) is the probability density function of $\tilde{y}$ [2][5]. In Formula (2.2), the same distribution can be expressed in either $\tilde{x}$ or $\tilde{y}$ or $\tilde{z}$ , which are different representations of the same underlying random variable. Using Formula (2.2), Formula (2.3) gives the $\rho(\tilde{y},\mu_{y},\sigma_{y})$ for $x^{c}$ if $\rho(\tilde{x},x,\sigma)$ is Gaussian.
$\rho(\tilde{z})d\tilde{z}=$ $\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y};$ (2.2)
$y=x^{c}:\;\;\;\rho(\tilde{y},\mu_{y},\sigma_{y})=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\sigma}N(\frac{\tilde{y}^{\frac{1}{c}}-\mu}{\sigma});$ (2.3)
Viewed in the $f^{-1}(\tilde{y})$ coordinate, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ is $\rho(\tilde{x},\mu,\sigma)$ modulated by $\frac{d\tilde{x}}{d\tilde{y}}=1/f^{(1)}_{x}$, in which $f^{(1)}_{x}$ is the first derivative of $f(x)$ to $x$. A distributional zero of the uncertainty distribution happens when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$, while a distributional pole happens when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$. Zeros and poles provides the strongest local modulation to $\rho(\tilde{x},\mu,\sigma)$:
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the result distribution is same as the original distribution, as $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5]. A linear transformation generates neither distributional zero nor distributional pole according to Formula (2.2).
@@
In both Figure 2.2 and 2.2, the result probability density function in the $\tilde{z}$ representation resembles the input distribution more when the mode of the distributions is further away from either distributional pole or distributional zero. Such resemblance allows generic characterization of the output.
@@ 2.3 Statistical Taylor Expansion
Formula (2.2) gives the uncertainty distribution of an analytic function. However, normal scientific and engineering calculations usually do not care about the result distribution, but just few statistics of the result, such as the result mean and deviation [2][3]. Such simple statistics is the result of statistical Taylor expansion.
Let $\overline{x}$ and $\delta x$ be the sample mean and deviation of a random variable, and let $\mu$ and $\sigma$ as the distribution mean and deviation of the random variable. An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines bounded momentum $\zeta(n,\kappa)$, in which $0<\varrho,\kappa$ are the bounding ranges. Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value. To achieve $\overline{f(x)}=f(x)$ when $\sigma=0$, $\zeta(n)$ is normalized by $\zeta(0)$ so that $\zeta(0)=1$.
@@
$f(x+\tilde{x})=$ $f(x+\tilde{z}\sigma)=\sum_{n=0}^{\infty}\frac{f^{(n)}_{x}}{n!}\tilde{z}^{n}\sigma^{n};$ (2.5)
@@
$=$ $\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.7)
@@
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
$\delta^{2}f(x,y)=$ $\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
@@ 2.4 Bounding Symmetry and Asymptote
A bounding $[x-\varrho\sigma,x+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. To achieve mean-preserving, $\kappa$ determines $\varrho$. With mean-preserving bounding, Formula (2.11) and (2.12) give the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
@@
$\delta^{2}(x\pm y)$ $=\zeta_{x}(2)(\sigma_{x})^{2}+\zeta_{y}(2)(\sigma_{y})^{2};$ (2.12)
@@
$\delta^{2}(xy)$ $=\zeta_{x}(2)(\sigma_{x})^{2}y^{2}+x^{2}\zeta_{y}(2)(\sigma_{y})^{2}+\zeta_{x}(2)(\sigma_{x})^{2}\zeta_{y}(2)(\sigma_{y})^{2};$ (2.14)
@@ 2.4.1 Uniform Input Uncertainty
For uniform distribution, the bounding range is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.15) gives $\zeta(2n)$ while $\zeta(2n+1)=0$.
@@ 2.4.2 Gaussian Input Uncertainty
If $\kappa=\infty$, $\zeta(2n)=(2n-1)!!$!, which will cause Formula (2.32) for $\log(x\pm\delta x)$ and Formula (2.35) for $(x\pm\delta x)^{c}$ to diverge. Also, as stated later, $\kappa$ is determine by the sample count $N$. When $\kappa$ has an upper bound, Formula (2.4) becomes Formula (2.16):
@@
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$! according to Formula (2.18). When $\kappa=5,n<5$, $\left|\eta(2n)/(2n-1)!!-1\right|<10^{-3}$.
@@ 2.4.3 An Input Uncertainty with Limited Range
Formula (2.20) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$. Formula (2.21) shows its mean-preserving equation. Formula (2.22) shows its bounded momentum. Formula (2.23) shows its asymptote.
$\rho(\tilde{x},\mu,\sigma)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;\mu=2\lambda;\;\;\;\sigma=\sqrt{2}\lambda;$ (2.20)
@@ 2.5 One Dimensional Examples
$\frac{\overline{e^{x}}}{e^{x}}$ $=\sum_{n=0}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.25)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.26)
@@
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}\sigma^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.28)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.29)
@@
The result variance in statistical Taylor expansion shows the nature of the calculation, such as $\sigma\rightarrow P(e^{x})$, $\sigma\rightarrow\delta\sin(x)$, $P(x)\rightarrow\delta\log(x)$, and $P(x)\rightarrow P(x^{c})$.